OpenStudy (anonymous):

factor a^4+a^6 can somebody help me with this

OpenStudy (anonymous):

factor maybe? start with \(a^4(1+a^2)\)

OpenStudy (anonymous):

1a^4+a^6?

OpenStudy (mathsolver):

@satellite73 is correct

OpenStudy (anonymous):

but i did that is it correct?

OpenStudy (mathsolver):

I'm afraid you have ended up where you stared.

OpenStudy (anonymous):

yeah but when u multiply what @satellite73 did thats what u get

OpenStudy (anonymous):

a^4(1+a^2)

OpenStudy (anonymous):

is that what yoou do can you please explain

OpenStudy (mathsolver):

first you must understand that a^6=a*a*a*a*a*a and a^4=a*a*a*a so if you factorise these out you get (a*a*a*a)*(1+(a*a)) and that can be simplified to a^4(1+a^2)

OpenStudy (anonymous):

a^4+a^6 =(a^4)(1+a^2)

OpenStudy (anonymous):

i made a mistake it was a^4+a^2 so you do that the same way like a*a*a*a+a*a*a

OpenStudy (anonymous):

2 would be hcm right

OpenStudy (anonymous):

a^4+a^2 =(a^2)(a^2 -1) =a^2((a-1)(a+1)) =a=0 or a=1 or a=-1

OpenStudy (anonymous):

it could be either one of those?

OpenStudy (cwrw238):

no - thers no minus in this a^2(a^2 + 1) are the factors - you can't factor any further

OpenStudy (anonymous):

thank you all for helping me

OpenStudy (anonymous):

no problem. :)

OpenStudy (anonymous):

i appreciate it :)